Circular No. 6181

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

PKS 1622–297 J. R. Mattox, University of Maryland and Universities Space Research

Association; S. Wagner, Landessternwarte Heidelberg; T. A. McGlynn, Compton Observatory Science Support Center; M. Malkan, University of California, Los Angeles; J. F. Schachter, Harvard-Smithsonian Center for Astrophysics, communicate: “The EGRET instrument aboard the Compton Observatory indicates that the flux of PKS 1622–297 at E > 100 MeV continues at a high level on June 25, the last day for which data have been analyzed. A GRO Target of Opportunity observation is scheduled to begin on June 30 to better observe this flare with EGRET, COMPTEL, and OSSE. The pointing is scheduled to continue until July 5. Observations at other wavelengths are of tremendous interest. High-temporal-density optical, infrared, and mm observations are expected to be of particular value.”

1991 JX A. W. Harris and G. Hahn, DLR Institute for Planetary Exploration,

Berlin; and J. Davies, Joint Astronomy Centre, Hilo, write: “Thermal infrared observations of near-Earth asteroid 1991 JX were made at Mauna Kea with the U.K. Infrared Telescope (UKIRT; + CGS3 spectrometer; range 10–20 μm) as part of the UKIRT service observing program on June 16, when Δ = 0.05 AU. The 10.6-μm flux was < 220 mJy. Assuming a visual albedo of 0.4 for a Vesta-like object (Hicks and Grundy, IAUC 6177), the 10.6-μm flux upper limit indicates a spherical equivalent diameter for 1991 JX of < 0.34 km (a lower albedo would imply an even smaller diameter). We note that an albedo of 0.4, combined with an absolute magnitude H = 19.13 (S. Mottola, private communication), leads to a diameter estimate of 0.32 km, which is in excellent agreement with the 10.6-μm observations.”

NOVA AQUILAE 1995 B. Skiff, Lowell Observatory, notes that examination of original Palo-

mar Sky Survey prints for a precursor star, at the position published on IAUC 6174, shows nothing; the nearest star (B ∼ 20, neutral color) lies ∼ 1′′.7 southeast at α = 19^{h}05^{m}26^{s}.73 (± 0′′.2), δ = −1^{o}42′04′′.2 ± 0′′.3 (equinox J2000.0, epoch 1950 July 19.3 UT). Skiff further notes that it is possible that proper motion shifted the nova a couple of arcsec over the past 45 years, so searches on more recent plates are warranted.

1995 June 29 Daniel W. E. Green